OpenStudy (anonymous):

Write an equation of a line that passes through (8, 5) and (11, 14)

OpenStudy (dls):

\[\Huge y-y_1=m(x-x_1)\] \[\Huge m=\frac{y_2-y_1}{x_2-x_1}\] x1=8 y1=5